OpenStudy (anonymous):

Using the graph of f(x) = log10x below, approximate the value of y in the equation 10y = 3.

OpenStudy (anonymous):

y ≈ 0.01 y ≈ 1.48 y ≈ −2.01 y ≈ 0.78

OpenStudy (anonymous):

@Hero @jim_thompson5910

jimthompson5910 (jim_thompson5910):

The function is \[\Large f(x) = \log_{10}(x)\] right?

OpenStudy (anonymous):

Yep

jimthompson5910 (jim_thompson5910):

Replace f(x) with y. Then convert to exponential form. \[\Large f(x) = \log_{10}(x)\] \[\Large y = \log_{10}(x)\] \[\Large 10^{y} = x\]

jimthompson5910 (jim_thompson5910):

Then on the right side, replace x with 3 to get \[\Large 10^{y} = 3\]

jimthompson5910 (jim_thompson5910):

After all that, convert back to logarithmic form \[\Large 10^{y} = 3\] \[\Large y = \log_{10}(3)\]

jimthompson5910 (jim_thompson5910):

So what you need to do is use the graph to find the y coordinate of the point on the curve when x = 3

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

what is that y value?

OpenStudy (anonymous):

0.48?

OpenStudy (anonymous):

That's not a choice though.

hero (hero):

Alternatively, another approach is to realize that \(10^0 = 1\) \(10^1 = 10\) So that means the value of y must be \(0 < y < 1\)

hero (hero):

From this information you can easily eliminate the middle two options.

hero (hero):

Actually, option 2 might be a typo

OpenStudy (anonymous):

Ok, what I was doing earlier was plugging the choices into log(10)y. The closest I got was 0.01 = y and that disproved 10^y=3.

OpenStudy (anonymous):

Option 2? What do you think is the alternative?

hero (hero):

\(y \approx 0.48\)

hero (hero):

Or the typo could be in option 4

hero (hero):

I don't know which one, but if I had to pick, I'd choose option 2

hero (hero):

Talk to your teacher about it.

OpenStudy (anonymous):

That would make a lot of sense. I copied the answers over here correctly, but perhaps my source/test has an error.

hero (hero):

Your test?

OpenStudy (anonymous):

Assignment. Any way, she is unreachable today, so I'll just wing it. Thank you for your help.